### v1

@@ Abstract
Epistemic uncertainty is crucial for safety-critical applications and out-of-distribution detection tasks. Yet, we uncover a paradoxical phenomenon in deep learning models: an epistemic uncertainty collapse as model complexity increases, challenging the assumption that larger models invariably offer better uncertainty quantification. We propose that this stems from implicit ensembling within large models. To support this hypothesis, we demonstrate epistemic uncertainty collapse empirically across various architectures, from explicit ensembles of ensembles and simple MLPs to state-of-the-art vision models including ResNets and Vision Transformers—for the latter, we examine implicit ensemble extraction and decompose larger models into diverse sub-models, recovering epistemic uncertainty. We provide theoretical justification for these phenomena and explore their implications for uncertainty estimation.
@@ 1 Introduction
Bayesian deep learning provides us with a principled framework for quantifying uncertainty in complex machine learning models (MacKay, 1992; Neal, 1994). A key concept in this framework is epistemic uncertainty, which represents a model’s uncertainty about its predictions due to limited knowledge or data (Smith & Gal, 2018; Der Kiureghian & Ditlevsen, 2009). This form of uncertainty is distinct from aleatoric uncertainty, which captures inherent noise or randomness in the data (Kendall & Gal, 2017). Epistemic uncertainty is particularly useful in safety-critical applications and decision-making scenarios, where understanding the limits of a model’s knowledge can prevent costly errors.
Epistemic uncertainty is usually quantified using the mutual information between the model’s predictions and its parameters (Houlsby et al., 2011; Gal et al., 2017). For deep ensembles (Lakshminarayanan et al., 2017), which consist of multiple independently trained neural networks, this amounts to the mutual information between the ensemble member index and the models’ predictions. This is equivalent to the disagreement among ensemble members on a particular input (Smith & Gal, 2018): if the models in the ensemble produce widely varying predictions for a given sample, this is indicative of high epistemic uncertainty.
Intuitively, one might expect that as deep learning models grow in size and complexity, their capacity for epistemic uncertainty would increase. As Fellaji & Pennerath (2024) argue, “the more parameters a model has, the more likely it is to fit the data in multiple ways. Put another way, the posterior and thus the posterior predictive will tend to be flatter, making the epistemic uncertainty grow,” which is aligned with the conventional understanding of model complexity and uncertainty.
However, our work provides evidence for a simple yet paradoxical phenomenon: when constructing higher-order ensembles, ensembles of ensembles, we observe an epistemic uncertainty collapse111We initially observed this behavior in Spring 2021, published informally at https://blackhc.notion.site/Ensemble-of-Ensembles-Epistemic-Uncertainty-for-OoD-4a8df73b0d8942c8872aab1848a4393b, and have since confirmed it in multiple independent experiments that we share here.. This collapse occurs because individual ensembles, given sufficient size and training, converge to similar predictive distributions, causing inter-ensemble disagreement to vanish as the ensemble size grows.
We hypothesize that, similar to ensembles of ensembles, implicit ensembling might occur within the layers of large over-parameterized neural networks, potentially leading to significant underestimation of epistemic uncertainty for traditional uncertainty estimators that rely on final logits. Hence, similar to deep ensembles that have been found to offer better calibration (Ovadia et al., 2019), implicit ensembling may explain why larger models also appear more calibrated (Tran et al., 2022). Recent work by Fellaji & Pennerath (2024) provides additional evidence that the epistemic uncertainty collapse can occur even in simple over-parameterized MLPs trained on standard benchmark datasets.
@@
A wide range of applications relies on accurate uncertainty estimates. These include active learning, where uncertainty guides data acquisition; anomaly detection, where uncertainty can signal out-of-distribution inputs; and safety-critical systems, where understanding model confidence is crucial for responsible deployment. The potential underestimation of epistemic uncertainty in these contexts could lead to overconfident predictions and compromised decision-making processes, particularly in large over-parameterized neural networks.
@@
In the remainder of this paper, we provide both theoretical justification and empirical evidence for the epistemic uncertainty collapse phenomenon using toy examples, and MLPs trained on MNIST. Vice-versa, we show that implicit ensemble extraction, which decomposes larger models into sub-models, effectively recovers hidden ensemble structure and epistemic uncertainty on simple MLPs and state-of-the-art vision models based on ResNets (He et al., 2015) and Vision Transformers (Dosovitskiy et al., 2021).
@@ 2 Theoretical Framework
Bayesian Model Average. The Bayesian Model Average (BMA) lies at the core of Bayesian deep learning. Let be the distribution over model parameters , given observed data . Then, the predictive distribution for a new input is:
Information-Theoretic Quantities. To formally quantify and differentiate between aleatoric and epistemic uncertainty, we can use an information-theoretic decomposition (Kendall & Gal, 2017). Let be the predicted output, and be the model parameters. We define:
@@ 2.1 Deep Ensembles
Deep ensembles, introduced by Lakshminarayanan et al. (2017), have emerged as a simple yet effective method for approximating Bayesian inference in deep learning (Wilson & Izmailov, 2020). They have been shown to offer superior performance compared to more complex Bayesian approximations on many tasks (Ovadia et al., 2019), including calibration measurements (Guo et al., 2017), out-of-distribution detection, and classification with rejection (Ashukha et al., 2020). The core idea is to train multiple neural networks independently using different random initializations.
The BMA of a deep ensemble of models is:
where represents the parameters of the -th model in the ensemble.
Given ensemble members, we can estimate the epistemic uncertainty via the mutual information between the predicted class and the model index (Beluch et al., 2018):
@@ 3 Epistemic Uncertainty Collapse for Ensembles of Ensembles
Let us formally investigate a paradoxical phenomenon: the collapse of epistemic uncertainty when constructing ensembles of ensembles. Consider a scenario where we construct a higher-order ensemble by creating multiple deep ensembles, each comprised of models. Intuitively, one might expect this approach to yield more robust uncertainty estimates, as it effectively increases the diversity of our model space. However, our analysis has a surprising outcome: as we increase the size of each individual ensemble, the disagreement between these ensembles—and consequently, our estimate of epistemic uncertainty—vanishes.
Let be a set of deep ensembles, each containing models:
where represents the parameters of the -th model in the -th ensemble. For a given input , the predictive distribution of the -th ensemble is:
where we have dropped conditioning on the data for brevity. Consequently, the predictive distribution of the ensemble of ensembles is equal to the BMA over all individual members:
By more canonically defining to depend on and as categorical random variables with uniform distribution, we can rephrase the epistemic uncertainty as:
Infinite Sub-Ensemble Size. How does the epistemic uncertainty change with increasing size of the sub-ensembles? For this, we note that if we let ,
independent of . Hence, thanks to the central limit theorem, we have:
Epistemic Uncertainty Decomposition. To better understand the relationship between the epistemic uncertainty of a single ensemble and that of an ensemble of ensembles, we can leverage the chain rule of the mutual information. Denoting the epistemic uncertainty of a single ensemble consisting of all models as and the epistemic uncertainty of the ensemble of ensembles as , we have:
This decomposition shows that the epistemic uncertainty of a single ensemble can be expressed as the sum of the epistemic uncertainty of an ensemble of ensembles and the expected epistemic uncertainty within each ensemble (conditioned on the ensemble index ). That is, the more epistemic uncertainty is captured by the sub-ensembles, the less epistemic uncertainty remains for the ensemble of ensembles.
Implications for Large Models and Their Ensembles. The epistemic uncertainty collapse we have derived for ensembles of ensembles could have significant implications for epistemic uncertainty quantification in large neural networks, particularly for foundation models: as models grow in size and complexity, they may exhibit implicit ensembling effects within their layers, leading to a reduction in epistemic uncertainty estimates when ensembling them or even when using individual models.
More concretely, large neural networks can be viewed as a hierarchical composition of implicit ensembles because each layer can be thought of as an ensemble of neurons, and successive layers of the network as a whole can then be considered an ensemble of ensembles. As the depth and width increase, we thus might observe the same epistemic uncertainty collapse as we have derived in ensembles of ensembles. Thus, larger models might not necessarily provide better uncertainty quantification.
This hypothesis about implicit ensembling in large neural networks set the stage for our empirical investigation.
@@ 4 Empirical Results
In the following section, we present a series of experiments that not only demonstrate the epistemic uncertainty collapse in explicit ensembles of ensembles but also find parallels in the behavior of wide neural networks, providing evidence for the hypothesized effects of implicit ensembling. Details on the models, training setup, datasets, and evaluation are provided in Appendix A and in Appendix B in the appendix.
@@
These results demonstrate the epistemic uncertainty collapse phenomenon: the systematic reduction of epistemic uncertainty, visible as reduced prediction variance and thus higher overall confidence, even in regions far from the training data. This toy example highlights how implicit ensembling within wide neural networks could lead to underestimating epistemic uncertainty, particularly in data-sparse regions.
@@
These results provide compelling evidence for the epistemic uncertainty collapse due to implicit ensembling in neural networks. As the width of the MLP increases, we observe a clear trend of decreasing mutual information across all datasets: MNIST (LeCun & Cortes, 1998), Dirty-MNIST (Mukhoti et al., 2023), and Fashion-MNIST (Xiao et al., 2017). In Figure 3(a), the effect is most pronounced for the in-distribution MNIST test data, but the collapse is also clearly visible for out-of-distribution data (Fashion-MNIST & SVHN test sets). The decrease in mutual information indicates a reduction in the model’s epistemic uncertainty as it grows larger, despite maintaining similar accuracy. In Figure 3(b), the mean difference in uncertainty scores between in-distribution (MNIST) and out-of-distribution (Fashion-MNIST) samples also decreases with increasing model width, further corroborating the collapse of epistemic uncertainty. However, the AUROC for OoD detection using different uncertainty metrics slightly improves as the model width increases, which is in line with the results by Fellaji & Pennerath (2024), who report a significant deterioration of OoD performance for such MLP ensembles trained on CIFAR-10 but not MNIST.
This shows that epistemic uncertainty collapse due to implicit ensembling might be a pervasive phenomenon in neural networks, even in relatively simple MLP models.
@@ 4.1 Implicit Ensemble Extraction
To further substantiate the hypothesis that implicit ensembling is the underlying mechanism driving epistemic uncertainty collapse, we demonstrate that it is possible to mitigate this collapse by decomposing larger models into constituent sub-models. This approach, which we term implicit ensemble extraction, allows us to recover the diversity of an ensemble from a single large model, effectively reversing the collapse of epistemic uncertainty. This not only offers insights into the nature of epistemic uncertainty collapse but also points towards potential practical strategies for improving model robustness and out-of-distribution detection.
@@
In Figure 4, we see the effectiveness of this implicit ensemble extraction technique. The results demonstrate that this decomposition can recover much of the epistemic uncertainty of an ensemble from a single model, providing support for our hypothesis about the mechanism underlying epistemic uncertainty collapse:
Effectiveness of Model Decomposition. The dashed lines in both subfigures of Figure 4 represent the extracted ensemble, as described in the previous section. Remarkably, this decomposed model nearly recovers the mutual information scores of a fully trained ensemble with the same width factor despite being based on the weights of a single model, and its OoD detection performance approaches that of a narrower model. Since we can successfully extract an ensemble, this result supports our hypothesis that implicit ensembling is indeed occurring within larger models and that we can unlock this ensemble structure through decomposition.
@@
Limitations. These exploratory results provide initial evidence that ensemble extraction technique can unlock meaningful epistemic uncertainty, which is useful for downstream tasks. This is aligned with the hypothesis of implicit ensembling. At the same time, we already see that mutual information is not always the best uncertainty metric, the comparative performance changes depending on the model temperature, and the pool size of the best implicit ensemble is not always the same for different metrics and model architectures. More results that show this are shown in Appendix D.

### v2

@@ Abstract
Epistemic uncertainty is crucial for safety-critical applications and data acquisition tasks. Yet, we find an important phenomenon in deep learning models: an epistemic uncertainty collapse as model complexity increases, challenging the assumption that larger models invariably offer better uncertainty quantification. We introduce implicit ensembling as a possible explanation for this phenomenon. To investigate this hypothesis, we provide theoretical analysis and experiments that demonstrate uncertainty collapse in explicit ensembles of ensembles and show experimental evidence of similar collapse in wider models across various architectures, from simple MLPs to state-of-the-art vision models including ResNets and Vision Transformers. We further develop implicit ensemble extraction techniques to decompose larger models into diverse sub-models, showing we can thus recover epistemic uncertainty. We explore the implications of these findings for uncertainty estimation.
@@ 1 Introduction
Bayesian deep learning provides us with a principled framework for quantifying uncertainty in complex machine learning models (MacKay, 1992; Neal, 1994). A wide range of applications relies on accurate uncertainty estimates. These include active learning (Settles, 2009; Osband et al., 2022), where uncertainty guides data acquisition; anomaly detection (Tagasovska & Lopez-Paz, 2019; Mukhoti et al., 2023), where uncertainty can signal out-of-distribution inputs; and safety-critical systems, where understanding a model’s confidence is crucial for responsible deployment (Band et al., 2021).
A key concept in this framework is epistemic uncertainty, which represents a model’s uncertainty about its predictions due to limited knowledge or data (Smith & Gal, 2018; Der Kiureghian & Ditlevsen, 2009). This form of uncertainty is distinct from aleatoric uncertainty, which captures inherent noise or randomness in the data (Kendall & Gal, 2017). Epistemic uncertainty is usually quantified using the mutual information between the model’s predictions and its parameters (Houlsby et al., 2011; Gal et al., 2017). For deep ensembles (Lakshminarayanan et al., 2017), which consist of multiple independently trained neural networks, this amounts to the mutual information between the ensemble member index and the predictions. This is equivalent to the disagreement among ensemble members on a particular input (Smith & Gal, 2018): if the models in the ensemble produce widely varying predictions, this is indicative of high epistemic uncertainty.
As deep learning models grow in size and complexity, the quality of their uncertainty estimates generally improves (Tran et al., 2022). However, our work provides evidence for a simple yet important phenomenon: when constructing higher-order ensembles, ensembles of ensembles, we observe an epistemic uncertainty collapse111We initially observed this behavior in 2021, published informally at https://blackhc.notion.site/Ensemble-of-Ensembles-Epistemic-Uncertainty-for-OoD-4a8df73b0d8942c8872aab1848a4393b, and have since confirmed it in multiple independent experiments that we share here.. This collapse occurs because individual ensembles, given sufficient size and training, converge to similar predictive distributions, causing inter-ensemble disagreement to vanish as the ensemble size grows.
We hypothesize that, similar to ensembles of ensembles, implicit ensembling might occur within large neural networks, potentially leading to significant underestimation of epistemic uncertainty for traditional uncertainty estimators that rely on the final logits. Hence, similar to deep ensembles that have been found to offer better calibration (Ovadia et al., 2019), implicit ensembling may explain why larger models also appear more calibrated (Tran et al., 2022). A recent preprint by Fellaji & Pennerath (2024) provides additional evidence of this phenomenon occurring even in simple over-parameterized MLPs trained on standard benchmark datasets but does not offer an explanation.
We introduce implicit ensembling as a possible explanation for this phenomenon. Our theoretical analysis demonstrates uncertainty collapse in ensembles of ensembles, and we present experimental evidence of similar collapse in wider models. Additionally, we show that our proposed implicit ensemble extraction technique can recover epistemic uncertainty from large models.
Our work is of interest because epistemic uncertainty estimation is crucial for safety-critical applications and decision-making scenarios (Wen et al., 2021), where understanding model limitations can prevent costly errors. While the potential underestimation of epistemic uncertainty in large over-parameterized neural networks could lead to overconfident predictions and compromised decisions, our proposed implicit ensemble extraction technique offers a promising approach to recover and properly quantify this uncertainty.
@@
In the remainder of this paper, we examine epistemic uncertainty collapse in ensembles of ensembles and in wider neural networks. We offer theoretical justification for the former and propose implicit ensembling as a potential mechanism in the latter. We provide empirical evidence of such a collapse in wider models through experiments with toy examples and MLPs trained on MNIST. Vice-versa, we show that implicit ensemble extraction methods, which decompose larger models into sub-models, recover hidden ensemble structure and epistemic uncertainty in both simple MLPs and state-of-the-art vision models based on ResNets (He et al., 2015) and Vision Transformers (Dosovitskiy et al., 2021).
@@ 2 Theoretical Framework
Bayesian Model Average. The Bayesian Model Average (BMA) usually lies at the core of Bayesian deep learning. Let be the distribution over model parameters . We are interested in the posterior distribution given some observed data , but this is not necessary for the rest of this work. Thus, let us use as the distribution of interest over model parameters: here, it will be the distribution of possible model parameters that we obtain by running a training procedure given the training data and different random initializations.
Let be the predicted output with likelihood . Then, the predictive distribution for a new input is:
Information-Theoretic Quantities. To formally quantify and differentiate between aleatoric and epistemic uncertainty, we can use an information-theoretic decomposition (Smith & Gal, 2018). Let be the predicted output, and be the model parameters. We define:
@@
While the choice of uncertainty measures remains an active research area (Hüllermeier & Waegeman, 2021; Wimmer et al., 2023; Bengs et al., 2022; Shen et al., 2024a), mutual information has strong theoretical and empirical support. The mutual information between predicted outputs and model parameters directly captures the core idea of epistemic uncertainty: how much additional data collection could reduce our predictive uncertainty. This interpretation is used in both active learning (Houlsby et al., 2011; Gal et al., 2017) and Bayesian optimal experimental design (Rainforth et al., 2024), where it is known as the expected information gain. The effectiveness of this approach requires only that the training procedure is approximately Bayesian in its treatment of new data. This property has been demonstrated across many machine learning methods (Mlodozeniec et al., 2024; Kirsch, 2023). Recent studies confirm the strong empirical performance of this mutual information compared to alternative uncertainty measures (Schweighofer et al., 2024). Other approaches based on variance decomposition via the law of total variance have been proposed (Kendall & Gal, 2017); we briefly compare these methods in Section A.1.
@@ 2.1 Deep Ensembles
Deep ensembles, introduced by Lakshminarayanan et al. (2017), have emerged as a simple yet effective method for approximating Bayesian inference in deep learning (Wilson & Izmailov, 2020). For a deeper discussion of the connection between deep ensembles and approximate Bayesian inference, see Section A.2. They have been shown to offer superior performance compared to more complex Bayesian approximations on many tasks (Ovadia et al., 2019), including calibration measurements (Guo et al., 2017), out-of-distribution detection, and classification with rejection (Ashukha et al., 2020). The core idea is to train multiple neural networks independently using different random initializations. The BMA of a deep ensemble of models is:
where are the parameters of the -th model in the ensemble, drawn from the distribution of possible trained models . Given uniform index distribution , we can estimate the epistemic uncertainty via the mutual information between the predicted class and the model index (Beluch et al., 2018):
This is just a Monte-Carlo approximation of the mutual information between the predicted class and the parameter distribution where the ensemble members are seen as samples from .
@@ 3 Epistemic Uncertainty Collapse for Ensembles of Ensembles
Let us formally investigate a paradoxical phenomenon: the collapse of epistemic uncertainty when constructing ensembles of ensembles. A standard deep ensemble comprises independently trained models that approximate a Bayesian Model Average (BMA). A different way to “scale up” deep ensemblesis to create multiple deep ensembles and then average their predictions. One might expect that this ensemble of ensembles would yield more robust epistemic uncertainty estimates. However, our analysis has the opposite outcome: as we increase the size of each individual ensemble, the disagreement between these ensembles—and consequently, our estimate of epistemic uncertainty—vanishes.
3.1 Setup and Predictive Distribution
Let be the number of sub-ensembles, and let each sub-ensemble contain models. Denote the parameters of the -th model in the -th sub-ensemble by . Each sub-ensemble provides a predictive distribution:
Once we collect all sub-ensembles into , the aggregated predictive distribution is
Viewed as a single (larger) ensemble, this model is a mixture over all members. Denote by the random variable for the choice of sub-ensemble, and by the random variable for the choice of model within that sub-ensemble. Then each is chosen with probability , yielding a Bayesian mixture over members. The empirical epistemic uncertainty within sub-ensemble is given by and the empirical epistemic uncertainty across sub-ensembles is given by .
3.2 Limit of Increasing Sub-Ensemble Size
A key observation is that growing , holding fixed, causes all sub-ensembles to converge to identical predictive distributions, so any disagreement across sub-ensembles vanishes. Formally, for ,
since the LHS is a Monte-Carlo estimator of the RHS term. This implies that sub-ensemble indices become irrelevant: all sub-ensembles yield identical outputs in the limit. Consequently, the epistemic uncertainty of the ensemble of ensembles—the mutual information between the ensemble index () and the predicted label —collapses to zero as the predictions of become independent of the ensemble index. In other words,
We call this epistemic uncertainty collapse.
See Section A.3 for a more detailed and formal argument.
3.3 Chain Rule Decomposition
To understand how the epistemic uncertainty of a single ensemble relates to that of an ensemble of ensembles, note that
In words, the epistemic uncertainty of a large ensemble can be broken into two parts: disagreement across sub-ensembles plus disagreement within each sub-ensemble . As , each sub-ensemble individually captures a near-complete BMA, so there is negligible disagreement across sub-ensembles.
4 Implicit Ensembling
Having established the epistemic uncertainty collapse for explicit ensembles of ensembles, this raises the question whether similar phenomena occur within individual large neural networks. We hypothesize that as neural networks increase in width, they might naturally exhibit ensembling internally, leading to reduced epistemic uncertainty estimates. This would have important implications for uncertainty quantification in large models, including modern foundation models.
Concretely, we propose that large neural networks inherently function as “implicit ensembles,” where operations like matrix multiplication (or average pooling) effectively average across many units similar to averaging predictions in explicit ensembles. Under this hypothesis, wider networks would exhibit stronger internal ensembling effects, and ensembles of such networks would behave like ensembles of ensembles. This suggests that the epistemic uncertainty collapse demonstrated in Section 3.2 would manifest as model width increases, potentially limiting uncertainty quantification benefits from simply scaling model size.
This implicit ensembling hypothesis guides our empirical investigations in the subsequent section. We examine whether epistemic uncertainty decreases as MLP width increases despite stable accuracy, and whether we can extract ensemble-like structures from single large models to recover the collapsed uncertainty. We also investigate average pooling in vision models as another potential form of implicit ensembling, where activations from different receptive fields are aggregated in ways that might contribute to uncertainty collapse.
4.1 Connection to Neural Tangent Kernel Theory
The hypothesized epistemic uncertainty collapse in larger models connects elegantly with Neural Tangent Kernel (NTK) theory (Jacot et al., 2018; Lee et al., 2019; Yang, 2019). NTK theory characterizes infinitely wide neural networks trained with gradient descent, showing that under appropriate conditions (initialization and learning rate), these networks converge to Gaussian Process predictors determined by the NTK. That is, for a fixed architecture, networks converge to identical predictive functions regardless of their random initialization in the infinite-width limit. Different initializations can follow different optimization paths but ultimately reach the same functional form. This has profound implications for uncertainty estimation.
In an ensemble of infinitely wide networks, each member converges to the same predictor function. The disagreement between ensemble members, which is the foundation of epistemic uncertainty estimation, therefore vanishes completely. Any uncertainty measure based on this disagreement, such as mutual information between predictions and ensemble membership, collapses to zero. This directly parallels our earlier findings in Section 3.2: as sub-ensemble size increases, each sub-ensemble’s predictive distribution approaches the common NTK-governed predictor, eliminating inter-ensemble disagreement.
Thus, NTK theory offers a theoretical framework that could be related to our proposed implicit ensembling hypothesis, though the connection remains speculative. As neural networks grow wider, their behavior increasingly approaches the NTK limit where all initializations converge to identical predictive distributions. We hypothesize that individual network components may still maintain considerable internal diversity, but this diversity might become progressively masked at the output level due to averaging effects in wider layers and pooling operations. If this hypothesis is correct, as width increases, such implicit averaging mechanisms would grow stronger, causing the network’s collective behavior to approach the NTK-governed predictor despite diverse internal representations. While NTK characterizes the infinite-width limit where uncertainty vanishes at the output, the implicit ensembling framework might help explain how this phenomenon occurs in practical finite-width networks and suggest how hidden uncertainty could potentially be recovered by accessing these diverse internal components. We present initial evidence for this through implicit ensemble extraction in Figure 4.
@@ 5 Empirical Results
In the following section, we present a series of experiments that demonstrate the epistemic uncertainty collapse in explicit ensembles of ensembles and show similar patterns in the behavior of wide neural networks. While these patterns are consistent with the implicit ensembling hypothesis we have proposed, we note that other explanations might also account for these observations. Details on the models, training setup, datasets, and evaluation are provided in Appendix B and in Appendix C in the appendix.
@@
These results demonstrate the epistemic uncertainty collapse phenomenon: the systematic reduction of epistemic uncertainty, visible as reduced prediction variance and thus higher overall confidence, even in regions far from the training data. This toy example highlights how the proposed implicit ensembling within wide neural networks could lead to underestimating epistemic uncertainty, particularly in data-sparse regions.
In Section A.4 in the appendix, we observe that epistemic uncertainty collapse also occurs in ensembles of random forests (Breiman, 2001), a fundamentally different model class from neural networks. As we discuss in Section 6, this distinguishes epistemic uncertainty collapse from phenomena like feature collapse (Van Amersfoort et al., 2020) and neural collapse (Papyan et al., 2020) that are specific to deep neural networks and their training dynamics.
@@
These results demonstrate epistemic uncertainty collapse in wider neural networks. This pattern is consistent with our implicit ensembling hypothesis, though other explanations may also be possible. As the width of the MLP increases, we observe a clear trend of decreasing mutual information across all datasets: MNIST (LeCun & Cortes, 1998), Dirty-MNIST (Mukhoti et al., 2023), and Fashion-MNIST (Xiao et al., 2017). In Figure 3(a), the effect is most pronounced for the in-distribution MNIST test data, but the collapse is also clearly visible for out-of-distribution data (Fashion-MNIST & SVHN test sets). The decrease in mutual information indicates a reduction in the model’s epistemic uncertainty as it grows larger, despite maintaining similar accuracy. In Figure 3(b), the mean difference in uncertainty scores between in-distribution (MNIST) and out-of-distribution (Fashion-MNIST) samples also decreases with increasing model width, further corroborating the collapse of epistemic uncertainty. However, the AUROC for OoD detection using different uncertainty metrics slightly improves as the model width increases, which is in line with the results by Fellaji & Pennerath (2024), who report a significant deterioration of OoD performance for such MLP ensembles trained on CIFAR-10 but not MNIST.
This shows that epistemic uncertainty collapse might be a pervasive phenomenon in neural networks, even in relatively simple MLP models.
@@ 5.1 Implicit Ensemble Extraction
To further investigate the hypothesis that implicit ensembling might contribute to epistemic uncertainty collapse, we demonstrate that it is possible to mitigate this collapse by decomposing larger models into constituent sub-models. This approach, which we term implicit ensemble extraction, allows us to recover the diversity of an ensemble from a single large model, potentially reversing the collapse of epistemic uncertainty. This not only offers insights into the nature of epistemic uncertainty collapse but also points towards potential practical strategies for improving model robustness and out-of-distribution detection.
@@
In Figure 4, we see the effectiveness of this implicit ensemble extraction technique. The results demonstrate that this decomposition can recover much of the epistemic uncertainty of an ensemble from a single model, providing support for our hypothesis about implicit ensembling:
Effectiveness of Model Decomposition. The dashed lines in both subfigures of Figure 4 represent the extracted ensemble, as described in the previous section. Remarkably, this decomposed model nearly recovers the mutual information scores of a fully trained ensemble with the same width factor despite being based on the weights of a single model, and its OoD detection performance approaches that of a narrower model. The fact that we can successfully extract an ensemble structure suggests that diverse functional components may exist within larger models, which is consistent with our implicit ensembling hypothesis.
@@
Limitations. These exploratory results provide initial evidence that ensemble extraction technique can unlock meaningful epistemic uncertainty, which is useful for downstream tasks. While this is consistent with the implicit ensembling hypothesis, it does not prove that implicit ensembling is the mechanism responsible for uncertainty collapse in wide networks. At the same time, we already see that mutual information is not always the best uncertainty metric, the comparative performance changes depending on the model temperature, and the pool size of the best implicit ensemble is not always the same for different metrics and model architectures. More results that show this are shown in Appendix E.